# Revisiting Experience Replayable Conditions

### v1

@@ Abstract
Experience replay (ER) used in (deep) reinforcement learning is considered to be applicable only to off-policy algorithms. However, there have been some cases in which ER has been applied for on-policy algorithms, suggesting that off-policyness might be a sufficient condition for applying ER. This paper reconsiders more strict “experience replayable conditions” (ERC) and proposes the way of modifying the existing algorithms to satisfy ERC. To this end, instability of policy improvements is assumed to be a key in ERC. The instability factors are revealed from the viewpoint of metric learning as i) repulsive forces from negative samples and ii) replays of inappropriate experiences. Accordingly, the corresponding stabilization tricks are derived. As a result, it is confirmed through numerical simulations that the proposed stabilization tricks make ER applicable to an advantage actor-critic, an on-policy algorithm. In addition, its learning performance is comparable to that of a soft actor-critic, a state-of-the-art off-policy algorithm.
@@ 1 Introduction
In the past decade, (deep) reinforcement learning (RL) [46] has made significant progress and has gained attention in various application fields, including game AI [35, 51], robot control [19, 53], autonomous driving [5, 7], and even finance [16]. RL agents can optimize their action policies in a trial-and-error manner, even in nonlinear and unknown environments. However, the trial-and-error process is computationally expensive, and the low sample efficiency of RL remains a challenge.
@@
To improve sample efficiency, many RL algorithms utilize experience replay (ER) [31]. ER stores the empirical data obtained during the trial-and-error process into a replay buffer instead of immediately streaming them. By randomly replaying the stored empirical data for learning, the impact of each empirical data for learning, i.e. sample efficiency, can be increased. Although ER is a simple technique, it is highly useful due to its compatibility with recent deep learning libraries and abundant computational resources. Various variants have been proposed as attention to practicality of ER has increased: e.g. prioritization of important empirical data with high replay frequency [40, 39]; and regularization of learning for inappropriate empirical data [34, 43]. However, it is claimed that ER is only applicable to off-policy algorithms [10]. For example, proximal policy optimization (PPO), an on-policy algorithm, learns after collecting a certain amount of empirical data but generally does not reuse them [42].
@@
On the other hand, a survey of past cases confirms the application of ER to on-policy algorithms. For example, previous studies [58, 2] reported that ER was successfully applied to state–action–reward–state–action (SARSA) [46], a typical on-policy algorithm  [^1] Bejjani et al [2] modified SARSA to deep Q-networks (DQN) [32], an off-policy algorithm, in subsequent updates [2].. This SARSA is also used to learn value functions in soft actor-critic (SAC) [14] and twin delayed deep deterministic policy gradient (TD3) [11], both of which are off-policy algorithms  [^2] Theoretically, their learning rules corresponds to the expected SARSA [49], which is an off-policy algorithm, but their implementations are with a rough Monte Carlo approximation and would lack rigor.. In addition, the author have successfully applied ER to PPO and its variants without any learning breakdowns [26].
@@
From the above suggestive cases, it is considered that being an off-policy algorithm corresponds to a sufficient condition for “experience replayable conditions” (ERC). Then, what is necessary and sufficient conditions for ERC? This study hypothesizes that i) policy improvement algorithms in RL have the corresponding sets of empirical data that can be acceptable for learning, and that ii) ER can be applied to them if only the empirical data belonging to the respective sets are replayed. With this hypothesis, one can expect that off-policy algorithms naturally satisfy ERC because the acceptable set for them coincides with the whole set of empirical data. As a remark, based on previous cases, it is assumeed that there is no ERC for learning value functions (this assumption will become clear in the numerical verification of this paper).
@@
To demonstrate the validity of this hypothesis, this paper first reveals the instability factors that destabilize learning by ER. Specifically, one can focus on the fact revealed in Kobayashi [24] that the standard policy improvement algorithms can be derived as a triplet loss in metric learning [50], from the viewpoint of control as inference (CaI) [29]. That is, the instability factors of triplet loss, i.e. i) repulsive forces from negative samples (so-called hard negative[6, 54] and ii) replays of inappropriate experiences (so-called distribution shift[41, 55], are inherited by the policy improvement algorithms. These instability factors might be easily apparent by ER, causing learning breakdowns.
Figure 1: Concept of two stabilization tricks for satisfying ERC
@@
To alleviate the identified instability factors, the two corresponding stabilization tricks, counteraction and mining, are proposed based on an experience discriminator. For the hypothesized ERC, i) the counteraction is responsible for expanding the acceptable set of empirical data, and ii) the mining is responsible for selectively replaying the empirical data belonging to it (see Fig. 1). Hence, ERC should hold by setting the appropriate counteraction and mining.
@@
With the proposed stabilization tricks, it is verified whether ER is applicable to an advantage actor-critic (A2C) [33], an on-policy algorithm, by satisfying ERC. First, a rough grid search of hyperparameters is performed to identify the range of counteraction and mining in which ERC holds. Second, ablation tests for the counteraction and mining with the found hyperparameters show the needs for both. Finally, the A2C with the proposed stabilization tricks is evaluated on more practical multi-dimensional tasks, suggesting the learning performance comparable to SAC, which is the state-of-the-art off-policy algorithm.
@@
    It is revealed that the reason why ER is limited to off-policy algorithms is due to the two instability factors associated with triplet loss hidden in the policy improvement algorithms.
@@
    The two stabilization tricks, the counteraction and mining, are developed to alleviate the instability factors respectively.
@@ 2.1 Basics of reinforcement learning
Let’s briefly introduce a problem statement for RL [46]. RL aims to maximize the sum of future rewards, so-called return, under Markov decision process (MDP). Mathematically, MDP is defined with the tuple $(\mathcal{S},\mathcal{A},p_{e},r)$: $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively; $p_{e}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ gives the stochastic state transition function; and $r:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{R}$ defines the reward function for the current situation with $\mathcal{R}\subset\mathbb{R}$ the set of rewards.
Under the above MDP, an agent encounters the current state $s_{t}\in\mathcal{S}$ of an (unknown) environment at the current time step $t\in\mathbb{N}$. The agent decides its action $a_{t}\in\mathcal{A}$ according to its learnable policy $\pi(a_{t}\mid s_{t}):\mathcal{S}\mapsto\mathcal{A}$. By interacting with the environment by $a_{t}$, the agent enters the next state $s_{t+1}=:s_{t}^{\prime}\in\mathcal{S}$ according to $p_{e}(s_{t}^{\prime}\mid s_{t},a_{t})$. At the same time, the reward $r_{t}=r(s_{t},a_{t})$ is obtained from the environment.
@@
where $\rho^{\pi}$ denotes the probability to generate the state-action trajectory, $\tau_{t}=[s_{t}^{\prime},a_{t+1},\ldots]$, under $\pi$ and $p_{e}$. Here, the expected return can be defined as the following learnable value functions.
@@ 2.2 RL algorithms with experience replay
ER stores the empirical data, $(s,a,s^{\prime},r,b)$ ($b=\pi(a\mid s)\in\mathbb{R}_{+}$ the policy likelihood to generate $a$, which is used for the proposed method), in a replay buffer $\mathcal{B}$ with finite size $|\mathcal{B}|$ in FIFO format [31]. For the sake of simplicity, this paper focuses on ER for learning algorithms with each single transition, although ER for long-term sequences has also been developed [17, 20]. That is, the loss function w.r.t. the learnable functions in RL (i.e. $\pi$, $V$, and/or $Q$ mainly) to be minimized can be computed for each empirical data. The following minimization problem is therefore given under ER.
@@
where $f$ indicates the function(s) to be optimized and $\ell_{f}^{\mathrm{alg}}$ denotes the algorithm-dependent loss function for $f$ (even if other functions are used for computing it, they are not optimized through minimizing it).
@@ 2.2.1 Soft actor-critic
Here, two algorithms used in this paper are introduced briefly. The first one is SAC [14], which is known as a representative off-policy algorithm, as the comparison. SAC learns two functions, $\pi$ and $Q$, through minimization of the following loss functions  [^3] The actual implementation has two $Q$ functions to conservatively compute the target value..
@@
As for the optimization of $Q$, the expected SARSA [49], which can be an off-policy algorithm, is employed. Indeed, the target value of $Q$ is computed with $\pi$, which is different from the policy used for obtaining the empirical data (so-called the behavior policy). In addition, the optimization of $\pi$ is off-policy as well since $Q$ to be minimized in $\ell_{\pi}$ can be dependent on arbitrary policies (as like Lillicrap et al [30]). In this way, SAC belongs to the off-policy algorithm in total. As a result, SAC can optimize $\pi$ and $Q$ with arbitrary empirical data, making it possible to use ER freely.
@@ 2.2.2 Advantage actor-critic
As $V$ necessarily depends on $\pi$ by definition, the above learning rule for $V$ requires $\pi$-dependent $r$ (and $s^{\prime}$). The learning rule for $\pi$ is also derived based on the policy gradient theorem, and again the dependence of $\pi$ is assumed for $V$. In this way, A2C belongs to the on-policy algorithm in total. As a result, A2C should not be theoretically able to learn with ER, which stores the empirical data independent of $\pi$. Note that the policy gradient algorithm is theoretically converted to the off-policy algorithm by utilizing importance sampling [8], but due to computational instability and distribution shift, various countermeasures must be implemented [52, 57, 9] (see Section 6.1 in details).
@@ 2.3 Metric learning with triplet loss
where $m\geq 0$ denotes the margin between the positive and negative clusters. The max operator is applied to prevent divergence into negative infinity. That is, the anchor data can be embedded near the positive data while being away from the negative data to some extent.
Figure 2: Desired distance relationship in metric learning with triplet loss
@@ 3.1 Policy improvements via control as inference
In this concept, an optimal variable $O=\{0,1\}$ is introduced. The probability mass functions, $p(O=1\mid s)$ and $p(O=1\mid s,a)$, can be defined with $V$ and $Q$, respectively.
@@
where $\beta>0$ denotes the inverse temperature parameter. For convenience, $C$, which is numerically unknown, is substracted to satisfy $V(s)-C\leq 0$ and $Q(s,a)-C\leq 0$. Note that since $O$ is binary, the probability of $O=0$ can be given as $p(O=0\mid s)=1-p_{V}(s)$ and $p(O=0\mid s,a)=1-p_{Q}(s,a)$.
@@
where $\mathrm{KL}(\cdot\mid\cdot)$ denotes Kullback-Leibler (KL) divergence. The gradient of these two terms w.r.t. $\pi$, $g_{\pi}$, is derived as follows:
@@
Here, by assuming $\beta\to\infty$, this gradient can be simplified. In addition, $Q-V=A$ the advantage function can be approximated by TD error $\delta$ (see Section 2.2.2).
It can be seen that the minimization of eq. (13) using the surrogated gradient is consistent with the minimization of the A2C loss function $\ell_{\pi}^{\mathrm{A2C}}$ by (stochastic) gradient descent. The original minimization problem can be interpreted as trying to move the anchor data $\pi$ closer to the positive data $\pi^{+}$ and away from the negative data $\pi^{-}$ by employing KL divergence as the distance function $d$  [^4] Although KL divergence does not actually satisfy the definition of distance, it is widely used in probability geometry because of iits distance-like property, which is non-negative and zero only when two probability distributions coincide.. Compared to eq. (9), the margin $m$ and the max operator are not used, but this is because $m=0$ and $\pi^{+,-}$ is centered on $b$, preventing the divergence to infinity.
@@ 3.2 Instability factors hidden in triplet loss
As the policy improvements correspond to minimizing triplet loss, then its characteristics during learning should also inherit that of minimizing triplet loss. Ideally, the anchor data should approach the positive data, resulting in $\pi\to\pi^{+}$. On the other hand, minimization of triplet loss is different from simple supervised learning, and several factors that destabilize learning have been reported. This study assumes that these instability factors become apparent in the policy improvements combined with ER. In this paper, therefore, the following two instability factors are raised as ones that needs to be addressed.
First, with inappropriate selection of anchor, positive, and negative data, $\mathrm{KL}(\pi\mid\pi^{+})>\mathrm{KL}(\pi\mid\pi^{-})$ might be caused. In this case, the repulsion from $\pi^{-}$ is stronger than the attraction to $\pi^{+}$ and the optimal solution cannot be found, updating $\pi$ by the divergent behavior. This is known as hard negative [6, 54]. To alleviate this instability factor, the exclusion of hard-negative triplets and/or the regularization that suppresses the repulsion would be required.
Second, from all the triplets that can be constructed, only a few can be used for optimization. In fact, $\pi^{+}$ and $\pi^{-}$ are linked via the behavior policy $b$ in the policy improvements, and no arbitrary triplet can be constructed. This causes distribution shift, inducing a large bias in learning [41, 55]. To alleviate this instability factor, it is desirable to eliminate triplets that are prone to bias and/or to regularize the distribution of selected triplets.

### v2

@@ Abstract
Experience replay (ER) used in (deep) reinforcement learning is considered to be applicable only to off-policy algorithms. However, there have been some cases in which ER has been applied for on-policy algorithms, suggesting that off-policyness might be a sufficient condition for applying ER. This paper reconsiders more strict “experience replayable conditions” (ERC) and proposes the way of modifying the existing algorithms to satisfy ERC. In light of this, it is postulated that the instability of policy improvements represents a pivotal factor in ERC. The instability factors are revealed from the viewpoint of metric learning as i) repulsive forces from negative samples and ii) replays of inappropriate experiences. Accordingly, the corresponding stabilization tricks are derived. As a result, it is confirmed through numerical simulations that the proposed stabilization tricks make ER applicable to an advantage actor-critic, an on-policy algorithm. Moreover, its learning performance is comparable to that of a soft actor-critic, a state-of-the-art off-policy algorithm.
@@ 1 Introduction
In the past decade, (deep) reinforcement learning (RL) [51] has made significant progress and has gained attention in various application fields, including game AI [39, 57], robot control [22, 60], autonomous driving [7, 10], and even finance [19]. RL agents can optimize their action policies for unknown environments in a trial-and-error manner. However, the trial-and-error process is computationally expensive, and the low sample efficiency of RL remains a challenge.
@@
To improve sample efficiency, many RL algorithms utilize experience replay (ER) [34]. ER stores the empirical data obtained during the trial-and-error process into a replay buffer instead of immediately streaming them. By randomly and repeatedly replaying the stored empirical data for learning, the impact of each empirical data for learning, i.e. sample efficiency, can be increased. Although ER is a simple technique, it is highly useful due to its compatibility with recent deep learning libraries and abundant computational resources. However, it is claimed that ER is only applicable to off-policy algorithms [13]. For example, proximal policy optimization (PPO), an on-policy algorithm, learns after collecting a certain amount of empirical data but generally does not reuse them [47].
@@
Nevertheless, a survey of past cases suggests the efficacy of ER to on-policy algorithms. For example, previous studies [65, 3] reported that ER was successfully applied to state–action–reward–state–action (SARSA) [51], a typical on-policy algorithm  [^1] Bejjani et al [3] modified SARSA to deep Q-networks (DQN) [36], an off-policy algorithm, in subsequent updates [4].. This SARSA is also used to learn value functions in soft actor-critic (SAC) [17] and twin delayed deep deterministic policy gradient (TD3) [14], both of which are off-policy algorithms  [^2] Theoretically, their learning rules corresponds to the expected SARSA [55], which is an off-policy algorithm, but their implementations are with a rough Monte Carlo approximation and would lack rigor.. In addition, the author have successfully applied ER to PPO and its variants without any learning breakdowns [28] (also see Appendix B).
@@
From the above suggestive cases, it is expected that being an off-policy algorithm corresponds to a sufficient condition for “experience replayable conditions” (ERC). Then, what is necessary and sufficient conditions for ERC? This study hypothesizes that i) policy improvement algorithms in RL have the corresponding sets of empirical data that can be acceptable for learning, and that ii) ER can be applied to them if only the empirical data belonging to the respective sets are replayed. With these hypotheses, one can expect that off-policy algorithms naturally satisfy ERC because the acceptable set for them coincides with the whole set of empirical data. As a remark, based on the past cases, it is assumeed that there is no ERC in the context of learning value functions (this assumption will become clear in the numerical verification of this paper).
@@
To demonstrate the validity of hypotheses, this paper first reveals the instability factors that destabilize learning by ER. Specifically, one can focus on the fact revealed in Kobayashi [26] that the standard policy improvement algorithms can be derived as a triplet loss in metric learning [56], from the viewpoint of control as inference (CaI) [32]. That is, the instability factors of triplet loss, i.e. i) repulsive forces from negative samples (so-called hard negative[8, 61] and ii) replays of inappropriate experiences (so-called distribution shift[45, 62], are inherited by the policy improvement algorithms. Since the way of sampling data makes the instability factors active, ER also inherit them by randomly selecting empirical data, causing learning breakdowns.
Figure 1: Concept of two stabilization tricks for satisfying ERC: Under the hypothesis that different algorithms have different sets of acceptable empirical data, the first trick, counteraction, expands that set, while the second trick, mining, selects the empirical data to be replayed.
@@
To alleviate the identified instability factors, the two corresponding stabilization tricks, a counteraction and a mining, are proposed respectively based on an experience discriminator. For the hypothesized ERC, i) the counteraction is responsible for expanding the acceptable set of empirical data, and ii) the mining is responsible for selectively replaying the empirical data belonging to it (see Fig. 1). Hence, ERC should hold by setting both of the counteraction and mining appropriately.
@@
With the proposed stabilization tricks, it is verified whether ER is applicable to an advantage actor-critic (A2C) [37], an on-policy algorithm, by satisfying ERC. First, a rough grid search of hyperparameters is performed to identify the range of counteraction and mining in which ERC holds. Second, ablation tests for the counteraction and mining with the found hyperparameters show the needs for both. Finally, the A2C with the proposed stabilization tricks is evaluated on more practical multi-dimensional tasks, suggesting the learning performance comparable to SAC [17], which is the state-of-the-art off-policy algorithm.
@@
    It is revealed that the reason why ER is limited to off-policy algorithms is due to the two instability factors associated with the triplet loss hidden in the policy improvement algorithms.
@@
    The two stabilization tricks, the counteraction and mining, are developed to alleviate the corresponding instability factors respectively.
@@ 2.1 Basics of reinforcement learning
Let’s briefly introduce a problem statement of RL [51]. RL aims to maximize the sum of future rewards, so-called return, under Markov decision process (MDP). Mathematically, MDP is defined with the tuple $(\mathcal{S},\mathcal{A},p_{e},r)$: $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces  [^3] Continuous spaces are assumed without loss of generality to match the experiments in this paper., respectively; $p_{e}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,\infty)$ gives the stochastic state transition function; and $r:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{R}$ defines the reward function for the current situation with $\mathcal{R}\subset\mathbb{R}$ the set of rewards.
Under the above MDP, an agent encounters the current state $s_{t}\in\mathcal{S}$ of an (unknown) environment at the current time step $t\in\mathbb{N}$. The agent (stochastically) decides its action $a_{t}\in\mathcal{A}$ according to its learnable policy $\pi(a_{t}\mid s_{t}):\mathcal{S}\times\mathcal{A}\mapsto[0,\infty)$. By interacting with the environment by $a_{t}$, the agent enters the next state $s_{t+1}=:s_{t}^{\prime}\in\mathcal{S}$ according to $p_{e}(s_{t}^{\prime}\mid s_{t},a_{t})$. At the same time, the reward $r_{t}=r(s_{t},a_{t})$ is obtained from the environment.
@@
where $\rho^{\pi}$ denotes the probability to generate the state-action trajectory, $\tau_{t}=[s_{t}^{\prime},a_{t+1},\ldots]$, with the combination of $\pi$ and $p_{e}$. Here, the expected return can be defined as the following learnable value functions.
@@ 2.2 RL algorithms with experience replay
ER stores the empirical data, $(s,a,s^{\prime},r,b)$ ($b=\pi(a\mid s)\in\mathbb{R}_{+}$ the policy likelihood to generate $a$, which is used for the proposed method), in a replay buffer $\mathcal{B}$ with finite size $|\mathcal{B}|$ [34]. For the sake of simplicity, this paper focuses on ER for learning algorithms with each single transition, although ER for long-term sequences has also been developed [20, 23]. That is, the loss function w.r.t. the learnable functions in RL (i.e. $\pi$, $V$, and/or $Q$ mainly) to be minimized is required to be computed only with each empirical data. The following minimization problem is therefore given under ER.
@@
where $f$ indicates the function(s) to be optimized and $\ell_{f}^{\mathrm{alg}}$ denotes the algorithm-dependent loss function for $f$ (even if other functions are used for computing it, they are not optimized through minimizing it). Although various variants have been proposed as the attention to practicality of ER has increased so far (e.g. prioritizing the important empirical data with high replay frequency [44, 43] and ignoring the inappropriate empirical data [38, 48]), this paper employs the most standard implementation with the FIFO-type replay buffer and uniform sampling of empirical data.
@@ 2.2.1 Soft actor-critic
Here, two algorithms mainly used in this paper are introduced briefly. The first one is SAC [17], which is known as a representative off-policy algorithm, as the comparison. SAC learns two functions, $\pi$ and $Q$, through minimization of the following loss functions  [^4] The actual implementation has two $Q$ functions to conservatively compute the target value..
@@
As for the optimization of $Q$, the expected SARSA [55], which can be an off-policy algorithm, is employed. Indeed, the target value of $Q$ is computed with $\pi$, which is different from the policy used when obtaining the empirical data (so-called the behavior policy). In addition, the optimization of $\pi$ is off-policy as well since $Q$ to be minimized in $\ell_{\pi}$ can be dependent on arbitrary policies (as like Lillicrap et al [33]). In this way, SAC belongs to the off-policy algorithm in total. As a result, SAC can optimize $\pi$ and $Q$ with arbitrary empirical data, making it possible to use ER freely.
@@ 2.2.2 Advantage actor-critic
As $V$ necessarily depends on $\pi$ by definition, the above learning rule for $V$ requires $\pi$-dependent $r$ (and $s^{\prime}$). The learning rule for $\pi$ is also derived based on the policy gradient theorem, and again the dependence of $\pi$ in $V$ is assumed. In this way, A2C belongs to the on-policy algorithm in total. As a result, A2C should not be theoretically able to learn with ER, which stores the empirical data independent of $\pi$. Note that the policy gradient algorithm can theoretically be converted to the off-policy algorithm by utilizing importance sampling [11], but due to computational instability and distribution shift, various countermeasures must be implemented [58, 64, 12] (see Section 6.1 in details).
@@ 2.3 Metric learning with triplet loss
where $m\geq 0$ denotes the margin between the positive and negative clusters. The max operator is employed to preclude the divergence into negative infinity. As a result, the anchor data can be embedded near the positive data while being away from the negative data to some extent.
Figure 2: Desired distance relationship in metric learning with triplet loss: The anchor data $x$ should be in the same cluster as the positive data $x^{+}$, and the negative data $x^{-}$ should be taken from a different cluster away from it.
@@ 3.1 Policy improvements via control as inference
In this concept, an optimal variable $O=\{0,1\}$ is introduced. The probability mass functions for it, $p(O=1\mid s)$ and $p(O=1\mid s,a)$, can be defined with $V$ and $Q$, respectively.
@@
where $\beta>0$ denotes the inverse temperature parameter. That is, if $\beta$ is small, these probabilities are likely to be near $1/2$, increasing ambiguity; and if $\beta$ is large, they often converge to 00 or $1$, making them deterministic. $C$ is an unknown value for convenience to satisfy $V(s)-C\leq 0$ and $Q(s,a)-C\leq 0$ so that the above equations satisfy the definition of probability and corresponds to the maximum value of the value function. Since numerical computation is impossible with $C$ remaining, $C$ must be excluded from the learning rule that is eventually derived. Note that since $O$ is binary, the probability of $O=0$ can be given as $p(O=0\mid s)=1-p_{V}(s)$ and $p(O=0\mid s,a)=1-p_{Q}(s,a)$.
@@
where $\mathrm{KL}(\cdot\mid\cdot)$ denotes Kullback-Leibler (KL) divergence between two probability distributions. The gradient of these two terms w.r.t. $\pi$, $g_{\pi}$, is derived as follows:
@@
Here, by assuming $\beta\to\infty$, this gradient can be simplified. In addition, $Q-V=A$ the advantage function can be approximated by TD error $\delta$ (see Section 2.2.2). As mentioned above, when $\beta$ is large, $p(O=1)$ deterministically takes 00 or $1$, and with $\beta\to\infty$, $O=1$ is not obtained unless $V=Q=C$ (i.e. the value function is maximized by the optimized policy, which is consistent with eq. (1)).
The minimization of eq. (13) using the surrogated gradient is consistent with the minimization of the loss function for A2C, $\ell_{\pi}^{\mathrm{A2C}}$, by (stochastic) gradient descent. The original minimization problem can be interpreted as trying to move the anchor data $\pi$ closer to the positive data $\pi^{+}$ and away from the negative data $\pi^{-}$ by employing KL divergence as the distance function $d$  [^5] Although KL divergence does not actually satisfy the definition of distance, it is widely used in probability geometry because of its distance-like property, which is non-negative and zero only when two probability distributions coincide.. Compared to eq. (9), the margin $m$ and the max operator are not used, but this is because $m=0$ and $\pi^{+,-}$ is centered on $b$, preventing the divergence to infinity.
@@ 3.2 Instability factors hidden in triplet loss
As the policy improvements correspond to minimizing triplet loss, then its characteristics during learning should also inherit that of minimizing triplet loss. Ideally, the anchor data should approach the positive data, resulting in $\pi\to\pi^{+}$. On the other hand, the minimization of triplet loss is different from simple supervised learning problems, and several factors that destabilize learning have been reported. These instability factors are influenced by the way triplets are selected from the dataset. In the policy improvements, therefore, they are activated by randomly sampling empirical data from ER (and using them for learning). Under this assumption, the following two instability factors are raised, along with their solution guidelines noted in the context of metric learning, which should be useful for RL combined with ER.
First, selecting the inappropriate anchor, positive, and negative data might cause $\mathrm{KL}(\pi\mid\pi^{+})>\mathrm{KL}(\pi\mid\pi^{-})$. In this case, the repulsion from $\pi^{-}$ is stronger than the attraction to $\pi^{+}$ and the optimal solution cannot be found, updating $\pi$ by the divergent behavior. This is known as hard negative [8, 61]. To alleviate this instability factor, the exclusion of hard-negative triplets and/or the regularization that suppresses the repulsion would be required.
Second, from all the triplets that can be constructed, only a few can be used for optimization. In fact, $\pi^{+}$ and $\pi^{-}$ are linked via the behavior policy $b$ in the policy improvements, and no arbitrary triplet can be constructed. This might cause distribution shift, inducing a large bias in learning [45, 62]. To alleviate this instability factor, it is desirable to eliminate triplets that are prone to bias and/or to regularize the distribution of selected triplets.